From mar@roe.ac.uk Fri Mar 23 14:44:08 2012 Date: Fri, 23 Mar 2012 15:43:57 +0000 From: Mike Read To: VDFS Science Archive Team -- Clive Davenhall , Eckhard Sutorius , Mark Holliman , Nicholas Cross , Nigel Hambly , Rob Blake , Ross Collins , Stelios Voutsinas Cc: CCs for VDFS -- Andrew Lawrence , Bob Mann , Jim Emerson , Keith Noddle , Lorenzo Rimoldini , Mike Irwin , Norman Gray , Stephen Warren , Tom Kerr , Tom Shanks Subject: WFAU VDFS Science Archives team meeting minutes, 23rd March 2012 Present: ETWS, ACD, MAR, MSH, RSC, STV Apologies: NJC, RGM, NCH, JPE, AL, KTN, NG, RPB, C37, 89X DoNM: 10am Friday April 6th 2012 in the VH (NCH in the chair) Actions discharged this time: ----------------------------- ACTION: ETWS to sort out common SQL function definition script(s) and ensure schema browser parser works correctly with those. Done, though work continues on revised schema browser parser. Actions partly discharged but continuing: ----------------------------------------- ACTION: RPB to roll out latest curation client / DB server share configuration for routine ops use. Continues; all servers now done, ETWS to modify ingester to work with new shares. ACTION: MAR to update VSA pages with details of public survey releases. Continues, waiting on further releases. Actions carried forward from 9/03/12 meeting --------------------------------------------- ACTION: RPB to set up WSA mirror/sync for flat file DBs using MS SQL replication. Continues. ACTION: ETWS, with RPB, to develop an automated helper script to keep release DB files links updated. Continues; this one still has low priority. ACTION: RPB to complete the LTO-5 tape backups of VVV and SSA DBs and SSS flat file datasets. Continues; RPB noted that this would now be carried out using LTO-5 as the new drive and tapes would be arriving soon. Of course with the advent of LTO-6 ... Specific points and new actions: -------------------------------- Project management: Nothing to report. WFCAM, VISTA and VST updates: No news to report this time. Comments and issues arising from CASU minutes: No new minutes this time. Networking: Up to date. ETWS had adapted the CU1 code to navigate the duplication of data on disks at CASU. WSA/VSA/OSA/GES Operations: OSA - all available files have been CU3'ed and metadata are available. GES - ACD reported that he and NCH had worked on a draft ICD between WFAU and CASU. This had been circulated to the interested parties but no feedback had been received to date. Hardware and Systems: MSH reported the arrival and status of three new machines. The two new curation servers (Ini and Neferefre) are up and running and ready for ETWS to do a software install. Both machines have 64 cores and 128 Gb RAM. The new DB server, Ramses13 (30 Tb), is re-building its disk system following the swapping in of 2Tb disks. Ramses13 is likely to be the new VVV public server. MSH will be installing a VM on one of the new curation machines to act as a test curation server. Both new curation machines will be on the SRIF. MSH also noted that we will soon be needing a new KVM switch and Infiniband switch. Survey Data Release: VISTA public release update: VVV: still processing synoptic products; ETA late April VMC: DR1 has been prepared. MAR reported that it had passed some quick sanity checks, eg. source table only contained data from the two required fields. VIDEO: The mosaics for the public release had been transferred. However the latest version of CU13 was causing problems. Some of the issues had been worked around but the mosaicing section needed looking at by NJC on his return. It was decided to try the version of CU13 that last worked on a VIDEO release and then revert to the current codebase for later CUs. VIKING: The CU19 release process was going slower than expected. For example the outgest of the multiframe table had taken over 8 hours. However it was still expected to finish within a few days. VHS: DR1 released. (The latest VHS proprietary release had fallen over when creating neighbours with 2XMM. It was thought this would be fixed by an update to the 2XMM schema.) Software: ETWS reported: -- Still rewriting the Browser parser for VISTA releases in the VSA. -- Unfortunately some of the last ingests of the OSA data fell over due to schema and code errors. All is fixed now and the flat file access can be soon made available. -- Started working on the code for ingests of the OSA catalogue data. Non-survey Data Release: Nothing to report this time. Astrogrid deployment & Data Analysis services: MSH was still in the process of migrating the Astrogrid services. Miscellaneous: Thanks to ACD, who is leaving us next week, for all his work on GES and ATLAS. -- Scanned by iCritical.